Find the cost of painting the outer surface of 24 oil containers at $2.50 per square meter, the dimensions of each container being 60 centimeters by 40 centimeters by 56 centimeters.
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

Find the cost of painting the outer surface of 24 oil containers at $2.50 per square meter, the dimensions of each container being 60 centimeters by 40 centimeters by 56 centimeters.

the dimensions of the container are 60 cms by 40 cms by 56 cms
= 0.6 meters by 0.4 meters by 0.56 meters
There are 8 faces in a cuboid 4 faces will have same area. The other four faces will have same arae.
let area of one face be 0.6 * 0.4 = =0.24 sq. meters
The other face will be 0.4* 0.56 = 0.22 sq. meters
The total surafce area of the can will be
4 * 0.24 + 4* 0.22 sq.. meters
0.96 + 0.88 sq. meters = 1.84
Number of containers = 24 , cost of painting per container = $2.50
Toal cost = 240 * 2.5 * 1.84 $ 1104

Answer by ikleyn(53996)

.
Find the cost of painting the outer surface of 24 oil containers at $2.50 per square meter,
the dimensions of each container being 60 centimeters by 40 centimeters by 56 centimeters.
~~~~~~~~~~~~~~~~~~~~~~~~~


        As I read the solution in the post by @mananth, I was shocked.
        He writes "There are 8 faces in a cuboid 4 faces will have same area. The other four faces will have same arae."

        In opposite, a cuboid has 6 faces, and opposite faces have equal areas.

        So, the @mananth solution is a fatal nonsense.

        I came to bring a correct solution.


If a rectangular box has dimensions L (the length), W (the width) and H (the height), then
its surface area is

    S = 2*(LW + LH + HL).


In our case, L = 60 cm = 0.6m;  W = 40 cm = 0.4 m;  H = 56 cm = 0.56 m,  and

    the surface area is  2*(0.6*0.4 + 0.6*0.56 + 0.40*0.56) = 2*0.8 = 1.6 m^2.


So, the cost of the painting is $2.50 times 1.6,  or 4 dollars.    ANSWER

Solved correctly.